Oriole Supply Company Income Statement for the Fiscal Year Ended June 30, 2017($ thousands) Net sales Cost of goods sold Selling and administrative expenses Nonrecurring expenses Earnings before interest, taxes, depreciation, and amortization (EBITDA) Depreciation Earnings before interest and taxes (EBIT) Interest expense Earnings before taxes (EBT) Taxes (35%) Net income $2,110,000 1,460,000 317,000 27,000 $306,000 113,000 193,000 118,000 $75,000 26,250 $48,750 Oriole Supply Company Balance Sheet as of June 30, 2017($ thousands) Assets: Liabilities and Equity: Cash and marketable $835,000 102,000 41,400 $978,400 1,165,000 $2,143,400 1,313,000 863,800 $2,176,800 $4,320,200 $417,000 Accounts payable securities Accounts receivable Inventory Other current assets 711,000 Notes payable 1,161,000 Accrued income taxes 42,200 Total current liabilities Total current assets $2,331,200 Long-term debt Net plant and equipment 1,989,000 Total liabilities Common stock Retained earnings Total common equity Total assets $4,320,200 Total liabilities and equity Refer to the balance sheet and income statement for Oriole Supply Company for the year ended June 30, 2017 (Round Days sales outstanding and Times-interest-earned to 1 decimal place, e.g 12.5, Total debt ratio and Debt-to-equity ratio to 3 decimal places, e.g. 12.557 and all other answers to 2 decimal places, e.g. 12.55. Use 365 days for calculation.) Calculate the following ratios: a. Inventory turnover b. Days sales outstanding e. Total debt ratio f. Debt-to-equity ratio g. Times-interest-earned h. Cash coverage days c. Total asset turnover d. Fixed asset turnover
Answer #1

a) Inventory Turnover = Cost of Goods Sold/Inventory

Inventory turnover = $1,460,000/$1,161,000 = 1.26

b) Days' sales outstanding = Account Receivable/Sales * 365

DSO = $711,000/$2,110,000 * 365 = 123.0 days

c) Total asset turnover = Sales/Total Assets

Total asset turnover = $2,110,000/$4,320,200 = 0.49

d) Fixed asset turnover = Sales/Fixed Assets

Fixed asset turnover = $2,110,000/$1,989,000 = 1.06

e) Total debt Ratio = Total Debt/Total Assets

Total debt ratio = (102,000 + 1,165,000)/1,267,000 = 0.293

f) Debt-to-Equity Ratio = Total Debt/Total Equity

Total debt ratio = (102,000 + 1,165,000)/2,176,800 = 0.582

g) Time interest earned = EBIT/Interest Expense

Times interest earned = $193,000/$118,000 = 1.6

h) Cash coverage ratio = EBITDA/Interest Expense

Times interest earned = $306,000/$118,000 = 2.59